For each precipitate (PbCrO4, Bi, Cu2[Fe(CN)6], CdS, HgCl2. Hg2Cl2, Hg, HgI), write a net ionic equation and note colors of products.
Answer #1

i) Pb + 2 + CrO4-2 = PbCrO4

Yellow color

ii) 3 Bi + 3 = 3 Bi

Bright white

iii) Cu + + Fe (CN) 6 + 2 = Cu2 [Fe (CN) 6)]

Brown color

iv) Cd + 2 + S-2 = CdS

Yellow color

v) Hg + 2 + 2 Cl- = HgCl2

White color

vi) 2 Hg + + Cl2 = Hg2Cl2

White color

vii) Hg + = Hg

Silver colour

viii) Hg + + I- = HgI

Yellow color

If you liked the answer, please rate it in a positive way, you would help me a lot, thank you.